# Enhancing the Clique Local Decoder to Correct Length-2 Space Errors in the Surface Code

### v1

@@ II-B Surface Codes
For example, Fig. 1 depicts a surface code layout with code distance $d=3$. Data qubits are shown as white ’D’ circles, while ancilla qubits for bit-flip and phase-flip error detection are marked as green ’X’ and blue ’Z’, respectively. A phase-flip (’Z’) error on the red-highlighted data qubit is detected by its diagonally adjacent ’X’ ancillae (shaded in yellow). Similarly, bit-flip (’X’) errors are identified by neighboring ’Z’ ancillae, and ’Y’ errors trigger both. Repeated entanglement of data and ancilla qubits via stabilizer circuits enables syndrome extraction for error detection.
@@ II-D Considerations for Cryogenic Quantum Systems
Some prior works have explored fully cryogenic QEC decoding to bypass I/O bottlenecks. For instance, NISQ+ [29] performs all decoding within the cryogenic domain, effectively alleviating bandwidth constraints. However, fully cryogenic decoders face severe limitations in chip area, thermal management, and computational complexity, creating another resource bottleneck. Similar challenges are observed in recent cryogenic decoders like QECOOL [48] and QULATIS [49], where resource constraints become even more pronounced as logical qubit counts and code distances increase—key requirements for fault-tolerant quantum computing.

### v2

@@ II-B Surface Codes
For example, Fig. 1 depicts a surface code layout with code distance $d=3$. Data qubits are shown as white ‘D’ circles, while ancilla qubits for bit-flip and phase-flip error detection are marked as green ‘X’ and blue ‘Z’, respectively. A phase-flip (‘Z’) error on the red-highlighted data qubit is detected by its diagonally adjacent ‘X’ ancillae (shaded in yellow). Similarly, bit-flip (‘X’) errors are identified by neighboring ‘Z’ ancillae, and ‘Y’ errors trigger both. Repeated entanglement of data and ancilla qubits via stabilizer circuits enables syndrome extraction for error detection.
@@ II-D Considerations for Cryogenic Quantum Systems
Some prior works have explored fully cryogenic QEC decoding to bypass I/O bottlenecks. For instance, NISQ+ [26] performs all decoding within the cryogenic domain, effectively alleviating bandwidth constraints. However, fully cryogenic decoders face severe limitations in chip area, thermal management, and computational complexity, creating another resource bottleneck. Similar challenges are observed in recent cryogenic decoders like QECOOL [44], where resource constraints become even more pronounced as logical qubit counts and code distances increase—key requirements for fault-tolerant quantum computing.